# Signed Higher-Order Interactions for Brain Disorder Diagnosis via Multi-Channel Transformers

### v3

@@ 3.2 Topological Higher-order and Lower-order organization of fMRI signals
Many studies have explored higher-order structures in brain networks by focusing on loops (1D cycles) formed by nodes and edges (Bian et al., 2024; Bhattacharya et al., 2025). Besides, the influence of signed HOIs was overlooked in these studies, which could offer a broader diagnostic view of neurological conditions. Our proposed strategy for representing signed HOIs allows researchers to explore more complex signed higher-order structures, such as quadriplet structures and voids (2D holes) formed by triangular faces. This advance is significant because it further refiness the HOIs by defining sign and links HOIs in the brain to higher-order organisations, and provides a higher-dimensional perspective for studying higher-order structures in the brain, potentially providing new insights into brain organisation and function.
3.2.1 Filtrate the Weighted Simplicial Complexes
To mine the topological higher-order and lower-order organization of fMRI signals, we employ Persistent Homology (Talesh Jafadideh and Mohammadzadeh Asl, 2022), a computational topology technique adept at analyzing high-dimensional datasets. The core concept involves constructing a filtration—a sequence of simplicial complexes that progressively approximate the original weighted simplicial complex with increasing precision:
In our study, we construct two distinct filtrations for the signed weighted simplicial complexes and . One filtration aims to extract all signed higher-order ”good” quadruplets structures from the signed weighted simplicial complex. These structures represent co-fluctuations that are not observable through lower-order interactions alone. The other filtration is designed to identify signed 2-dimensional holes, which reflect the intrinsic shape of the data from a higher-dimensional perspective. Additionally, we also extract all lower-order edge structures from the weighted simplicial complex .
The first filtration. Sort the weights of edges, triangles, and quadruplets in ascending order for the signed weighted simplicial complexes. The parameter scans this sequence, effectively tracking the current weight as we progress through the sorted list. At each step , we identify all quadruplets that satisfy the simplicial downward closure condition, i.e., for each quadruplet , the following conditions hold:
where , , and denote the weights of the quadruplet, triangle, and edge, respectively. These quadruplets, along with their corresponding weights, are added to the signed list of ”good” quadruplets and . Additionally, all edges are added to the list of edges .
The second filtration. Sort the weights of edges, triangles, and quadruplets in descending order for the signed weighted simplicial complexes. At each step , we include all triangles and quadruplets that satisfy the simplicial upward closure condition. These triangles and quadruplets and all edges, along with their corresponding weights, are added to the signed list of persistent homology and .
@@ 3.2.1 Filtrate the Weighted Simplicial Complexes
Persistent homology examines how higher-order topological features evolve through the filtration , offering a measure of their robustness across scales. By focusing on 2D voids within the homology group , we track these higher-order organization. Specifically, we apply Persistent Homology to the signed list of persistent homology and to generate two two-dimensional persistence diagrams. In each diagram, each point signifies a void that emerges during the filtration process. The persistence quantifies the void’s lifespan, indicating its importance.
3.2.2 Homological scaffold and edge projection
@@
The homological scaffold reveals the roles of different links in shaping the system’s homological properties. A larger total persistence for a link indicates its function as a locally strong bridge within the space of coherent and decoherent co-fluctuations. To analyze the information from the signed list of good quadruplets and at the edge levels, we use edge projections (Barrat et al., 2004). For each edge , we assign a weight equal to the average sum of the weights of the quadruplets defined by that edge. The feature matrices of 2D voids using homological scaffold and the feature matrices of quadruplet signatures using edge projections are illustrated in Fig. 3.
At each time point , five distinct weighted networks are constructed: one derived from edge structures representing lower-order interactions, two from signed good tetrahedral structures signifying HOIs, and another two from signed 2D holes illustrating the data’s intrinsic geometric shape from a higher-dimensional perspective. By averaging across the entire time period, the resulting networks exhibit enhanced stability and mitigate the impact of fMRI noise.
@@ 3.3.1 Signed higher-order Features Decoupling Mechanism
Higher-order features that are positive are related to patterns of simultaneous activation in multiple brain regions at a given moment, while higher-order features that are negative are related to patterns of collective inhibition in these regions at the current moment. These features can offer valuable information about complex coordination and communication within the brain. We further explore the influence of signed higher-order features using higher-order features decoupling mechanism. Specifically, for each type of higher-order feature’s two signed weighted adjacency matrix (), we extract its upper triangular part and flatten it to construct a feature vector . We generate two new higher-order topological features by adaptively learning weights to aggregate positive and negative information:
where denotes a two-layer multi-layer perceptron (MLP), respectively, and denotes features of quadruplet structures and that aggregate positive and negative information, respectively. The attention score emphasizes discriminative signed features while suppressing redundant signed information.
@@ 3.3.2 Multi-Head Self-Attention Module of Multiple Channels
The feature matrices extracted above actually represents the low-order and higher-order topological information associated with each brain region. Simply concatenating the low-order and higher-order information of each brain region as the input to the model would overlook the heterogeneity of the related information. Therefore, inspired by Brain network Transformer (Kan et al., 2022b), we leverage a -layer non-linear mapping module, namely Multi-Head Self-Attention (MHSA), to generate more expressive node features using the low-order and higher-order topological features of multiple channels . For each layer , the output of each channel is obtained by
where , is the concatenation operator, is the number of heads, is the layer index, , , , are learnable model parameters of each channel, and is the first dimension of .
@@ 3.3.3 Orthonormal Clustering Readout of Multiple Channels
Research has shown that low-level and high-level feature patterns in the brain often exhibit distinct functional modular structures, thereby supporting the richest functional interactions (Wang et al., 2019). Therefore, to leverages the properties of brain networks in different patterns that nodes in the same functional modules tend to have similar behaviors and clustered representations, we design a orthonormal clustering readout of multiple channels inspired by the reference (Kan et al., 2022b) to take advantage of the modular-level similarities between ROIs in different patterns of brain networks, where nodes in different patterns are assigned softly to well-chosen clusters with an unsupervised process.
Formally, given cluster centers where each center has dimensions in given channel i, , a Softmax projection operator calculates the probability of assigning node to cluster :
where denotes the inner product, and is the learned set of node embeddings from the last layer of the MHSA module in given channel i. Using this soft assignment (where is the number of nodes), the graph-level embedding is obtained by:
@@ 3.3.4 Attention-Guided Feature fusion Mechanism for Classification
For embedding at the graph level of each channel obtained by flattening (), we ignite an attention-guided feature fusion mechanism to emphasize discriminative features while suppressing redundant information. The formula can be described as:
where denotes sigmoid function, and is element-wise multiplication, and are learnable model parameters. Subsequently, a three-layer multi-layer perceptron (MLP) is employed using as input for the prediction head. Finally, the cross-entropy loss function is used as the loss function.
@@ 4.1.1 Datasets and Preprocessing
To better evaluate the proposed method, we employ four fMRI datasets with varying data sizes. The class distribution of brain network datasets is shown in Table 1. 1) ADNI dataset (Dadi et al., 2019). The primary goal of ADNI has been to test whether serial magnetic resonance imaging (MRI), positron emission tomography (PET), other biological markers, and clinical and neuropsychological assessment can be combined to measure the progression of mild cognitive impairment (MCI) and early Alzheimer’s disease (AD). The raw images used in this paper were obtained from the Alzheimer’s Disease Neuroimaging Initiative (ADNI) database, which contains 90 Alzheimer’s disease (AD), 96 normal controls (NC) and 76 mild cognitive impairment (MCI). 2) TaoWu dataset (Xu et al., 2023). The TaoWu datasets, released by ICI (Xu et al., 2023), is among the earliest image datasets made available for Parkinson’s research, which contains 20 Parkinson’s disease (PD), 20 normal controls (NC). 3) PPMI dataset (Xu et al., 2023). The Parkinson’s Progression Markers Initiative (PPMI) is a comprehensive study aiming to identify biological markers associated with Parkinson’s risk, onset, and progression. The raw images used in this article were obtained from the Parkinson’s progression markers initiative (PPMI) database, which contains 53 Parkinson’s disease (PD) and 53 prodromal. 4) ABIDE dataset (Craddock et al., 2013). The Autism Brain Imaging Data Exchange (ABIDE) initiative supports the research on Autism Spectrum Disorder (ASD) by aggregating functional brain imaging data from laboratories worldwide. The raw images used in this article were obtained from the the Autism Brain Imaging Data Exchange (ABIDE) initiative database, which contains 488 Autism spectrum disorder (ASD) patients and 537 normal controls (NC) . We followed in resting-state fMRI data preprocessing (Esteban et al., 2019), which involves a number of pipeline steps, including motion correction, realigning, field unwarping, normalization, bias field correction, and brain extraction. AAL atlas (Tzourio-Mazoyer et al., 2002) is applied to the preprocessed images to parcellate the brain into 90 ROIs, and the BOLD time series of every ROI were obtained by averaging the time series of all voxels inside the ROI.
@@ 4.1.2 Methods for comparison
The selected baselines correspond to five categories. The first category is traditional machine learning models by the scikit-learn library (Pedregosa et al., 2011) using edge features, including Multilayer Perceptron (MLP) , Support Vector Machine Classifier (SVM), Logistic Regression, and Random Forest. The second category consists of GNN-based models that learn lower-order topological features from a two-way network. These models include GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GroupINN (Yan et al., 2019), BrainGNN (Li et al., 2021), FBNetGen (Kan et al., 2022a), BPI-GNN (Zheng et al., 2024). The third category consists of Transformer-based models that utilize the Transformer architecture to perceive the entire brain picture, including Graph Transformer (Ying et al., 2021), Brain Network Transformer (Kan et al., 2022b), Tsen (Hu et al., 2023), Long-range Brain Transformer (Yu et al., 2024). The fourth category is higher-order methods based on hypergraph, including HGCN (Hao et al., 2023), HGAT (Wang et al., 2022). The fifth category is persistent homology (PH)-based models, including Brain-HORS (Santoro et al., 2024), PH-MCI (Bhattacharya et al., 2025), ATPGCN (Bian et al., 2024).
@@ 4.1.3 Implementation Details
For all experiments,we evaluated the performance in terms of diagnosis accuracy, recall, precision, and F1 score. For dataset partitioning, we adopt an 8:1:1 ratio for training, validation and testing for the ADNI dataset and ABIDE dataset, while using a 3:1:1 ratio for the TaoWu and PPMI datasets. When evaluating model performance, we implement 10-fold cross-validation for the ADNI dataset and ABIDE dataset while 5-fold cross-validation for the others. All models are trained using the Adam optimiser (Kingma and Ba, 2014) with an initial learning rate of and a weight decay of , in conjunction with the ReduceLROnPlateau scheduler (Paszke et al., 2017). The batch size is set to 16 for the ADNI dataset and ABIDE dataset while 4 for others, with a maximum number of epochs of 100. After each epoch, the model performance is evaluated using the validation set. We keep the model with the highest F1 score for testing. All experiments are run within the PyTorch framework (Paszke et al., 2017). More details can be described as follows:
@@
For traditional machine learning models, the correlation matrix is computed by Pearson correlation for all time series, and its flattened upper triangular matrix is taken as a feature vector. A MLP with three fully connected layers is applied for classification. A SVM with a radial basis function kernel is applied for classification. A LogisticRegression with solver=lbfgs is applied for classification. A RandomForest with n_estimators=300 is applied for classification.
@@
For GNN-based models and Transformer-based models, the correlation matrix is calculated using Pearson correlation for all time series. On this basis, the node features in the graph are the columns of the corresponding Pearson correlation matrix. The adjacency matrix in the graph is obtained by thresholding (0.5) the correlation matrix. Generally, two graph convolutional layers and an average pooling layer are used for classification.
@@
For hypergraph-based higher-order methods, specifically, for HGCN, a graph is first constructed through the correlation matrix, where the correlation of each edge of the graph is greater than 0.7. Each node constructs 2-hyperedges, 3-hyperedges, and 4-hyperedges by selecting its k-hop neighbors in the graph. For HGAT, the feature of the central node corresponding to each hyperedge is used as the hyperedge feature. When aggregating information, the attention coefficient between the hyperedge and the node is calculated for classification. Note that more HGNN-based methods applying to brain data have been proposed recently. However, the code of these models is notpublicly available, and they require significant engineering which is hard to reproduce fairly based solely on the papers.
@@
For the persistent homology (PH)-based models, including PH-MCI, ATPGCN. We use the persistent topological features of connected components and loops ) extracting by Ripserer tool. For Brain-HORS, following the method, we extract the matrix of violating triangles, the matrix of one-dimensional loops and the matrix of lower-order edge features, then flatten them and feed them into the SVM for classification.

### v4

@@ 4.3 Topological Lower-Order and Higher-Order Organizations of fMRI Signals
Many studies have explored higher-order structures in brain networks by focusing on loops (1D cycles) formed by nodes and edges (Bian et al., 2024; Bhattacharya et al., 2025), with limited attention paid to more complex higher-order organizational patterns. Besides, the influence of signed HOIs was overlooked in these studies, which could offer a broader diagnostic view of neurological conditions. The new definition of weighted simplicial complexes allows researchers to explore more complex signed higher-order structures, such as quadriplet structures and voids (2D holes) formed by triangular faces. This advance is significant because it further refiness the HOIs by defining sign and links HOIs in the brain to higher-order organisations, and provides a higher-dimensional perspective for studying higher-order structures in the brain.
Capture lower-order edges: To explore brain activity patterns for diagnosis from a global perspective, we extract all the 1-simlices in the monotonic weighted simplicial complex and the monotonic weighted simplicial complex repesenting the lower-order organizations in brain and add them into a list .
In this paper, we focus on the impact of signed HOIs, so we do not further explore the signs of lower-order edges. We construct a weighted matrix from the list and assign a weight equal to the weight of the edge for each element in matrix.
Capture higher-order good quadruplets: To mine the topological higher-order of fMRI signals, we first extract all 3-simplices from the monotonic weighted simplicial complexes and and add them into two lists, respectively. These structures represent co-fluctuations that are not observable through lower-order interactions alone and represent extremely good higher-order organizations.
To quantify these good quadruplets at the edge levels from the lists, we use edge projections (Barrat et al., 2004). For each edge , we assign a weight equal to the average sum of the weights of the quadruplets defined by that edge. The feature matrices of quadruplet signatures using edge projections are illustrated in Figure 4.
Capture higher-order 2D voids: Similarly, we can extract all 3-simplices from the monotonic weighted simplicial complexes and to extract that represent extremely bad higher-order organizations. To some extent, these higher-order organizations can represent the 2-dimensional voids in the brain that are extracted by Persistent Homology, which reflect the intrinsic shape of the data from a higher-dimensional perspective. Therefore, we employ Persistent Homology (Talesh Jafadideh and Mohammadzadeh Asl, 2022) to extract 2-dimensional voids, which not only capture extremely bad higher-order organization but also help characterize the intrinsic geometric organization of neural activity from the perspective of a higher-dimensional manifold. More details about Persistent Homology therory and 2D void, please refer to A. The core concept involves constructing a filtration—a sequence of simplicial complexes that progressively approximate the original weighted simplicial complex with increasing precision:
Finally, the is equal to the monotonic weighted simplicial complexes or . Persistent homology examines how higher-order topological features evolve through the filtration , offering a measure of their robustness across scales. It is worth noting that calculating Persistent Homology requires defining the distance among multiple nodes in k-simplex, which can be easily achieved by computing (for ), where represents the weight of the k-simplex.
@@
By focusing on 2D voids within the homology group , we track these higher-order organization. Specifically, we apply Persistent Homology to the monotonic weighted simplicial complexes and to generate two two-dimensional persistence diagrams. In each diagram, each point signifies a void that emerges during the filtration process. The persistence quantifies the void’s lifespan, indicating its importance.
@@
The homological scaffold reveals the roles of different links in shaping the system’s homological properties. A larger total persistence for a link indicates its function as a locally strong bridge within the space of coherent and decoherent co-fluctuations. The feature matrices of 2D voids using homological scaffold are illustrated in Figure 4.
At each time point , five distinct weighted networks are constructed: one derived from edge structures representing lower-order interactions, two from signed good quadruplet structures signifying HOIs, and another two from signed 2D holes illustrating the data’s intrinsic geometric shape from a higher-dimensional perspective. By averaging across the entire time period, the resulting networks exhibit enhanced stability and mitigate the impact of fMRI noise.
@@ 4.4.1 Signed Higher-Order Features Decoupling Mechanism
Higher-order features that are positive are related to patterns of simultaneous activation in multiple brain regions at a given moment, while higher-order features that are negative are related to patterns of collective inhibition in these regions at the current moment. These features can offer valuable information about complex coordination and communication within the brain. We further explore the influence of signed higher-order features using higher-order features decoupling mechanism. Specifically, for each type of higher-order feature, we consider two signed weighted matrices (). For each of these matrices, we extract its upper triangular part and flatten it to construct a feature vector . We generate two new higher-order topological features by adaptively learning weights to aggregate positive and negative information:
where denote two separate two-layer multi-layer perceptrons (MLPs), and denote, respectively, features of quadruplet structures and features of 2D voids that aggregate positive and negative information, respectively. The attention scores emphasize discriminative signed features while suppressing redundant signed information.
@@ 4.4.2 Multi-Head Self-Attention Module of Multiple Channels
The feature matrices extracted above actually represent the low-order and higher-order topological information associated with each brain region separately. Simply feeding the concatenated low-order and higher-order information into the model would overlook the heterogeneity of the related information. Therefore, inspired by Brain network Transformer (Kan et al., 2022b), we leverage an -layer non-linear mapping module, namely Multi-Head Self-Attention (MHSA), to generate more expressive node features using the low-order and higher-order topological features of multiple channels . Specifically, for each channel , the output of each layer is obtained by
where is the concatenation operator, is the number of heads, is the head index, is the layer index, , , , are learnable model parameters of each channel i, and is the first dimension of .
@@ 4.4.3 Orthonormal Clustering Readout of Multiple Channels
Research has shown that low-level and high-level feature patterns in the brain often exhibit distinct functional modular structures, thereby supporting the richest functional interactions (Wang et al., 2019). Therefore, to leverage the inherent properties of brain networks, where nodes within the same functional module tend to exhibit similar behaviors and form clustered representations, we design an orthonormal clustering readout of multiple channels inspired by the reference (Kan et al., 2022b). This module is intended to capture modular-level similarities between ROIs across different brain network patterns, where nodes are softly assigned to well-defined clusters through an unsupervised process.
Formally, given cluster centers where each center has dimensions in a given channel i, , a Softmax projection operator calculates the probability of assigning node to cluster :
where denotes the inner product, denotes the cluster index, is the learned embeddings of node j from the last L-th layer of the MHSA module in given channel i, and denotes the embeddings of cluster k in given channel i. Using this soft assignment (where is the number of nodes), the graph-level embedding is obtained by:
@@ 4.4.4 Attention-Guided Feature Fusion Mechanism for Classification
For embedding at the graph level of each channel i obtained by flattening (), we propose an attention-guided feature fusion mechanism to emphasize discriminative features while suppressing redundant information. The formula can be described as:
where denotes sigmoid function, is element-wise multiplication, and , are learnable model parameters. Subsequently, a three-layer multi-layer perceptron (MLP) is employed using as input for the prediction head. Finally, the cross-entropy loss function is used as the loss function.
@@ 5.1.1 Datasets and Preprocessing
To comprehensively evaluate the proposed method, we employ four fMRI datasets with varying sample sizes. The class distribution of brain network datasets is presented in Table 1. 1) ADNI dataset (Dadi et al., 2019): The Alzheimer’s Disease Neuroimaging Initiative (ADNI) primarily aims to investigate whether the combination of serial magnetic resonance imaging (MRI), positron emission tomography (PET), other biological markers, along with clinical and neuropsychological assessments, can effectively measure the progression of mild cognitive impairment (MCI) and early Alzheimer’s disease (AD). The dataset comprises 90 Alzheimer’s disease (AD) patients, 96 cognitively normal (CN) subjects, and 76 individuals with mild cognitive impairment (MCI). 2) TaoWu dataset (Xu et al., 2023): Released by the International Conference on Imaging (ICI) (Xu et al., 2023), the TaoWu dataset represents one of the earliest neuroimaging collections for Parkinson’s disease research, containing 20 Parkinson’s disease (PD) patients and 20 normal controls (NC). 3) PPMI dataset (Xu et al., 2023): The Parkinson’s Progression Markers Initiative (PPMI) is a landmark longitudinal study designed to identify biomarkers of Parkinson’s disease risk, onset, and progression. Our study utilizes data from 53 Parkinson’s disease (PD) patients and 53 prodromal cases obtained from the PPMI database. 4) ABIDE dataset (Craddock et al., 2013): The Autism Brain Imaging Data Exchange (ABIDE) initiative facilitates Autism Spectrum Disorder (ASD) research by aggregating functional brain imaging data from multiple international sites. Our analysis includes 488 ASD patients and 537 normal controls (NC) from the ABIDE database.
All resting-state fMRI data were preprocessed following the pipeline described in (Esteban et al., 2019), which includes motion correction, realignment, field unwarping, normalization, bias field correction, and brain extraction. The preprocessed images were parcellated into 90 regions of interest (ROIs) using the Automated Anatomical Labeling (AAL) atlas (Tzourio-Mazoyer et al., 2002). The blood-oxygen-level-dependent (BOLD) time series for each ROI were obtained by averaging the time series of all constituent voxels.
@@ 5.1.2 Methods for Comparison
The selected baselines correspond to five categories. The first category consists of traditional machine learning models by the scikit-learn library (Pedregosa et al., 2011) using edge features, including Multilayer Perceptron (MLP), Support Vector Machine Classifier (SVM), Logistic Regression, and Random Forest. The second category consists of GNN-based models that learn lower-order topological features from a two-way network. These models include GCN (Kipf and Welling, 2017), GraphSAGE (Hamilton et al., 2017), GAT (Veličković et al., 2018), GroupINN (Yan et al., 2019), BrainGNN (Li et al., 2021), FBNetGen (Kan et al., 2022a), BPI-GNN (Zheng et al., 2024). The third category consists of Transformer-based models that utilize the Transformer architecture to perceive the entire brain network, including Graph Transformer (Ying et al., 2021), Brain Network Transformer (Kan et al., 2022b), Tsen (Hu et al., 2023), Long-range Brain Transformer (Yu et al., 2024). The fourth category consists of higher-order methods based on hypergraph, including HGCN (Hao et al., 2023), HGAT (Wang et al., 2022). The fifth category consists of Persistent Homology (PH)-based models, including Brain-HORS (Santoro et al., 2024), PH-MCI (Bhattacharya et al., 2025), ATPGCN (Bian et al., 2024).
@@ 5.1.3 Implementation Details
For all experiments, we evaluated the performance in terms of diagnosis accuracy, recall, precision, and F1 score. For dataset partitioning, we adopt an 8:1:1 ratio for training, validation and testing for the ADNI dataset and ABIDE dataset, while using a 3:1:1 ratio for the TaoWu and PPMI datasets. When evaluating model performance, we implement 10-fold cross-validation for the ADNI dataset and ABIDE dataset while 5-fold cross-validation for the others. All models are trained using the Adam optimiser (Kingma and Ba, 2014) with an initial learning rate of and a weight decay of , in conjunction with the ReduceLROnPlateau scheduler (Paszke et al., 2017). The batch size is set to 16 for the ADNI dataset and ABIDE dataset while 4 for others, with a maximum number of epochs of 100. After each epoch, the model performance is evaluated using the validation set. We keep the model with the highest F1 score for testing. All experiments are run within the PyTorch framework (Paszke et al., 2017). Further implementation details are summarized below:
@@
For traditional ML models, feature vectors were derived from the flattened upper triangle of Pearson correlation matrices computed from BOLD time series. An MLP with three fully connected layers is applied for classification. An SVM with a radial basis function kernel is applied for classification. A LogisticRegression with solver=lbfgs is applied for classification. A RandomForest with n_estimators=300 is applied for classification.
@@
For GNN-based models and Transformer-based models, the correlation matrix is calculated using Pearson correlation for all time series. Node features are derived from the columns of the Pearson correlation matrix. The adjacency matrix in the graph is obtained by thresholding (0.5) the correlation matrix. Generally, two graph convolutional layers and an average pooling layer are used for classification.
@@
For hypergraph-based higher-order methods, specifically, for HGCN, a graph is first constructed through the correlation matrix, where the correlation of each edge of the graph is greater than 0.7. Each node constructs 2-hyperedges, 3-hyperedges, and 4-hyperedges by selecting its k-hop neighbors in the graph. For HGAT, the feature of the central node corresponding to each hyperedge is used as the hyperedge feature. During information aggregation, attention coefficients between hyperedges and nodes are computed for classification. Note that more HGNN-based methods applying to brain data have been proposed recently. However, the code of these models is not publicly available, and they require significant engineering which is difficult to reproduce faithfully based solely on the papers.
@@
For PH-based models (e.g., PH-MCI and ATPGCN), we use the persistent topological features of connected components and loops extracted by Ripserer tool. For Brain-HORS, following the method, we extract the matrix of violating triangles, the matrix of one-dimensional loops and the matrix of lower-order edge features, which are then flattened and classified using an SVM.